given this equation y-3=3(x+1), what is the equation in standard form of a perpendicular line that passes thru (5,1) and what is the x intercept of the perpendicular line
Answer by lwsshak3(11628)

given this equation y-3=3(x+1), what is the equation in standard form of a perpendicular line that passes thru (5,1) and what is the x intercept of the perpendicular line.
***
standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
equation of given line:
y-3=3(x+1)
y-3=3x+3
y=3x+6
slope=3
..
slope of line perpendicular to given line=-1/3 (negative reciprocal)
equation of perpendicular line: y=(-1/3)x+b=-x/3+b
solve for b using coordinates of given point (5,1)
1=-1*5/3+b
b=8/3
equation: y=-x/3+8/3
find x-intercept:
set y=0
-x/3+8/3=0
x/3=8/3
x=8
x-intercept=8